# Fully distributed consensus control for stochastic multi-agent systems under undirected and directed topologies

### 14 Mar 2026 v1

@@ IV-A Fully distributed consensus control for stochastic MASs with the directed topology
$\begin{split}\mathrm{E}\int_{0}^{\infty}\eta_{1}\varepsilon\xi^{T}(s)[\tilde{S}(\tilde{C}(s)\tilde{\Sigma}(s))\otimes P]\xi(s)ds<\infty.\end{split}$
Combining the above inequality with Barbalat’s lemma [36], we have $\lim_{t\to\infty}\mathrm{E}\xi^{T}(t)[\eta_{1}\varepsilon\tilde{S}(\tilde{C}(t)\tilde{\Sigma}(t))\otimes P]\xi(t)=0$, that is, $\lim_{t\to\infty}\mathrm{E}|(\tilde{C}(t)\tilde{\Sigma}(t))|\|\xi(t)\|^{2}=0$. Note that $\xi(t)=(\mathcal{L}\otimes I_{n})\vartheta(t)$ and $\lim_{|\vartheta|\to\infty}\inf_{t\geq 0}V(\vartheta_{t},t)=\infty$. Using the stochastic LaSalle lemma [37, Corollary 3.1], we can get $\lim_{t\to\infty}\xi(t)=0$ a.s. According to the definition of $\xi(t)$, the m.s. and a.s. consensus can be solved. ∎
@@
In fact, the function $\Sigma_{i}$ admits other design forms. Especially, when we choose $\Sigma_{i}(t)=k_{1}(k_{2}+\sigma_{i}(t))^{\mu}$ with $\mu>1,k_{1},k_{2}\geq 1$ and $c_{i}(t)\geq 1$, $\Sigma_{i}(t)\geq 1$. By selecting the Lyapunov function $\tilde{V}(\vartheta_{t},t)=V(\vartheta_{t},t)$, it can be deduced that $\mathfrak{L}\tilde{V}(\vartheta_{t},t)=\sum_{i=1}^{M}r_{i}[\eta_{1}^{l}\Sigma_{i}(t)\sigma_{i}(t)+2\eta_{2}^{l}c_{i}(t)-2\eta_{2}^{l}\psi_{i}]\dot{c}_{i}(t)+\eta_{1}^{l}\sum_{i=1}^{M}r_{i}c_{i}(t)\Sigma_{i}(t)\mathfrak{L}\sigma_{i}(t)+\sum_{i=M+1}^{N}s_{i-M}[\eta_{1}^{f}\Sigma_{i}(t)\sigma_{i}(t)+2\eta_{2}^{f}c_{i}(t)-2\eta_{2}^{f}\psi_{i}]\dot{c}_{i}(t)+\eta_{1}^{f}\sum_{i=M+1}^{N}s_{i-M}c_{i}(t)\Sigma_{i}(t)\mathfrak{L}\sigma_{i}(t)$. By the similar steps of the above proof of Theorem IV.1, the m.s. and a.s. consensus of (1) can be solved by the protocol (2) with $\Sigma_{i}(t)=k_{1}(k_{2}+\sigma_{i}(t))^{\mu}$, for $\mu>1,k_{1},k_{2}\geq 1$ and $c_{i}(0)\geq 1$, $\Sigma_{i}(0)\geq 1$.
@@
Note that the Lyapunov functional $V(\vartheta_{t},t)$ in Theorem IV.1 satisfies $V(\vartheta_{t},t)\leq\sum_{i=1}^{M}r_{i}[\eta_{1}^{l}k_{1}(2k_{2}c_{i}(t)+\sigma_{i}(t))\sigma_{i}(t)+\eta_{2}^{l}(c_{i}(t)-\psi_{i})^{2}]+\sum_{i=M+1}^{N}s_{i-M}[\eta_{1}^{f}k_{1}(2k_{2}c_{i}(t)+\sigma_{i}(t))\sigma_{i}(t)+\eta_{2}^{f}(c_{i}(t)-\psi_{i})^{2}]$ with $\mu=1$. Then, we choose Lyapunov functional
$\begin{split}V_{1}&(\vartheta_{t},t)\\ =&\sum_{i=1}^{M}\!r_{i}\!\Big[\eta_{1}^{l}k_{1}(2k_{2}c_{i}(t)\!+\!\sigma_{i}(t))\sigma_{i}(t)\!+\!\eta_{2}^{l}(c_{i}(t)\!-\!\psi_{i})^{2}\Big]\\ &\!\!+\!\!\!\!\sum_{i=M+1}^{N}\!\!\!\!\!s_{i-M}\!\Big[\eta_{1}^{f}k_{1}(2k_{2}c_{i}(t)\!+\!\sigma_{i}(t))\sigma_{i}(t)\!+\!\eta_{2}^{f}(c_{i}(t)\!-\!\psi_{i})^{2}\Big],\end{split}$
where $\eta_{1}^{l},\eta_{2}^{l},\eta_{1}^{f},\eta_{2}^{f}>0$, $\psi_{i}$ are positive constants to be determined. According to the functional $\mathrm{It\hat{o}}$ formula, it can be deduced that $dV_{1}(\vartheta_{t},t)$ has the following form
$\begin{split}dV_{1}&(\vartheta_{t},t)\!=\!\mathfrak{L}V_{1}(\vartheta_{t},t)dt+\frac{\partial V_{1}(\vartheta_{t},t)}{\partial\vartheta}(I_{N}\!\otimes\!C)\vartheta(t)dw(t),\end{split}$
where $\mathfrak{L}V_{1}(\vartheta_{t},t)=\sum_{i=1}^{M}r_{i}\big[(\eta_{k_{2}}^{l}\sigma_{i}(t)+2\eta_{2}^{l}c_{i}(t)-2\eta_{2}^{l}\psi_{i})\dot{c}_{i}(t)\big]+\sum_{i=1}^{M}r_{i}\big[\big(\eta_{k_{2}}^{l}c_{i}(t)+\eta_{k_{1}}^{l}\sigma_{i}(t)\big)\mathfrak{L}\sigma_{i}(t)\big]+\sum_{i=M+1}^{N}s_{i-M}\big[(\eta_{k_{2}}^{f}\sigma_{i}(t)+2\eta_{2}^{f}c_{i}(t)-2\eta_{2}^{f}\psi_{i})\dot{c}_{i}(t)\big]+\sum_{i=M+1}^{N}s_{i-M}\big[\big(\eta_{k_{2}}^{f}c_{i}(t)+\eta_{k_{1}}^{f}\sigma_{i}(t)\big)\mathfrak{L}\sigma_{i}(t)\big]$ with $\eta_{k_{1}}^{l}=2\eta_{1}^{l}k_{1}$, $\eta_{k_{2}}^{l}=2\eta_{1}^{l}k_{1}k_{2}$, $\eta_{k_{1}}^{f}=2\eta_{1}^{f}k_{1}$, and $\eta_{k_{2}}^{f}=2\eta_{1}^{f}k_{1}k_{2}$. Denote $\tilde{\psi}=\mathrm{diag}\{\psi_{1},...,\psi_{N}\}$. According to Lemma IV.1, $\tilde{\psi}$ can be partitioned into $\tilde{\psi}_{l}=\mathrm{diag}\{\psi_{1},...,\psi_{M}\}$ and $\tilde{\psi}_{f}=\mathrm{diag}\{\psi_{M+1},...,\psi_{N}\}$. Let $\mathcal{K}=-B^{T}P$ and $\Gamma=PBB^{T}P$. Note that $\sum_{i=1}^{M}\sigma_{i}(t)=\xi_{l}^{T}(t)(I_{M}\otimes P)\xi_{l}(t)$ and $\xi_{l}(t)=(\mathcal{L}_{11}\otimes I_{n})\vartheta_{l}(t)$. By the definition of $\sigma_{i}(t)$, (24), (25) and using Lemma IV.1 and IV.2, we have
@@
Denote $\Theta^{l}_{t}=\eta_{k_{2}}^{l}\tilde{C}_{l}(t)+\eta_{k_{1}}^{l}\tilde{\sigma}_{l}(t)$ and $\Theta^{f}_{t}=\eta_{k_{2}}^{f}\tilde{C}_{f}(t)+\eta_{k_{1}}^{f}\!\tilde{\sigma}_{f}(t)$. Substituting the above inequalities into $\mathfrak{L}V_{1}(\vartheta_{t},t)$ yields
$\begin{split}\mathfrak{L}&V_{1}(\vartheta_{t},t)\\ \leq&\xi_{l}^{T}\!(t)[R\Theta^{l}_{t}\!\otimes\!(A^{T}\!P\!+\!PA\!+\!C^{T}\!PC)]\xi_{l}(t)\\ &+\xi_{f}^{T}\!(t)[S\Theta^{f}_{t}\!\otimes\!(A^{T}\!P\!+\!PA\!+\!C^{T}\!PC)]\xi_{f}(t)\\ &-2\tilde{\eta}_{l}\xi_{l}^{T}\!(t)\big[(k_{1}k_{2}\tilde{C}_{l}(t)\!+\!k_{1}\tilde{\sigma}_{l}(t))^{2}\!\otimes\!\Gamma\big]\xi_{l}(t)\\ &-2\tilde{\eta}_{f}\xi_{f}^{T}(t)\big[(k_{1}k_{2}\tilde{C}_{f}(t)+k_{1}\tilde{\sigma}_{f}(t))^{2}\otimes\Gamma\big]\xi_{f}(t)\\ &+\xi_{l}^{T}(t)[R(\eta_{k_{2}}^{l}\tilde{\sigma}_{l}(t)\!+\!2\eta_{2}^{l}\tilde{C}_{l}(t)\!-\!2\eta_{2}^{l}\tilde{\psi}_{l})\!\otimes\!\Gamma]\xi_{l}(t)\\ &+\xi_{f}^{T}(t)[S(\eta_{k_{2}}^{f}\tilde{\sigma}_{f}(t)\!+\!2\eta_{2}^{f}\tilde{C}_{f}(t)\!-\!2\eta_{2}^{f}\tilde{\psi}_{f})\otimes\Gamma]\xi_{l}(t),\end{split}$
@@
$\begin{split}\mathfrak{L}&V_{1}(\vartheta_{t},t)\\ \leq&\xi_{l}^{T}\!(t)[R\Theta^{l}_{t}\!\otimes\!(A^{T}\!P\!+\!PA\!+\!C^{T}\!PC\!+\!k_{2}\Gamma)]\xi_{l}(t)\\ &\!\!\!\!+\!\xi_{f}^{T}\!(t)[S\Theta^{f}_{t}\!\otimes\!(A^{T}\!P\!+\!PA\!+\!C^{T}\!PC\!+\!k_{2}\Gamma)]\xi_{f}(t)\\ &\!\!\!\!-\!\xi_{l}^{T}\!(t)\big[(2\tilde{\eta}_{l}(k_{1}k_{2}\tilde{C}_{l}(t)\!+\!k_{1}\tilde{\sigma}_{l}(t))^{2}\!+\!2\eta_{2}^{l}\tilde{\psi}_{l}R)\!\otimes\!\Gamma\big]\xi_{l}(t)\\ &\!\!\!\!-\!\xi_{f}^{T}(t)\big[(2\tilde{\eta}_{f}(k_{1}k_{2}\tilde{C}_{f}(t)\!+\!k_{1}\tilde{\sigma}_{f}(t))^{2}\!+\!2\eta_{2}^{f}\tilde{\psi}_{f}S)\otimes\Gamma\big]\xi_{f}(t)\\ \leq&\xi_{l}^{T}\!(t)[R\Theta^{l}_{t}\!\otimes\!(A^{T}\!P\!+\!PA\!+\!C^{T}\!PC\!+\!k_{2}\Gamma)]\xi_{l}(t)\\ &\!\!\!\!+\!\xi_{f}^{T}\!(t)[S\Theta^{f}_{t}\!\otimes\!(A^{T}\!P\!+\!PA\!+\!C^{T}\!PC\!+\!k_{2}\Gamma)]\xi_{f}(t)\\ &\!\!\!\!-\!2\xi_{l}^{T}\!(t)\Big[\sqrt{\tilde{\eta}_{l}\eta_{2}^{l}\tilde{\psi}_{l}R}\Big(k_{1}k_{2}\tilde{C}_{l}(t)\!+\!k_{1}\tilde{\sigma}_{l}(t)\Big)\!\otimes\!\Gamma\big]\xi_{l}(t)\\ &\!\!\!\!-\!2\xi_{f}^{T}(t)\big[\sqrt{\tilde{\eta}_{f}\eta_{2}^{f}\tilde{\psi}_{f}S}\Big(k_{1}k_{2}\tilde{C}_{f}(t)+k_{1}\tilde{\sigma}_{f}(t)\Big)\otimes\Gamma\Big]\xi_{f}(t).\end{split}$
@@
$\begin{split}\mathfrak{L}V_{1}(\vartheta_{t},t)\leq&-\!\xi_{l}^{T}\!(t)[R\Theta^{l}_{t}\otimes I_{n}]\xi_{l}(t)\!-\!\xi_{f}^{T}\!(t)[S\Theta^{f}_{t}\otimes I_{n}]\xi_{f}(t)\\ \leq&-\!\varepsilon\xi_{l}^{T}\!(t)\Big[R\Big(\eta_{k_{2}}^{l}\tilde{C}_{l}(t)\!+\!\frac{\eta_{k_{1}}^{l}}{2}\tilde{\sigma}_{l}(t)\Big)\otimes P\Big]\xi_{l}(t)\\ &-\!\varepsilon\xi_{f}^{T}\!(t)\Big[S\Big(\eta_{k_{2}}^{f}\tilde{C}_{f}(t)\!+\!\frac{\eta_{k_{1}}^{f}}{2}\tilde{\sigma}_{f}(t)\Big)\otimes P\Big]\xi_{f}(t)\\ \leq&-\varepsilon V_{1}(\vartheta_{t},t)+\varphi(t),\end{split}$
@@
Theorem IV.1 and IV.2 designed fully distributed protocols to solve m.s. and a.s. consensus for stochastic MASs under directed graphs. Different from the existing adaptive protocol designs [24, 26] for stochastic MASs with undirected graphs, Theorem IV.1 and IV.2 considered the case of directed graphs and proposed a fully distributed control protocol that can be applied to solve the m.s. and a.s. consensus under directed graphs.
Furthermore, two time-varying gains $c_{i}(t)$ and $\Sigma_{i}(t)$ are incorporated in the adaptive strategy (2). These two terms serve distinct functions in facilitating stochastic consensus: $c_{i}(t)$ is utilized to estimate global information, while $\Sigma_{i}(t)$ addresses the asymmetrical interactions characteristic of directed graphs. Additionally, owing to the asymmetry inherent in the Laplacian matrix of a directed graph, we have adopted a Lyapunov function that diverges from that used in the case of undirected graphs, and utilized the properties of M-matrix to address this issue.
@@
In fact, the parameter selection criteria based on the Riccati equation remain fundamentally similar to the existing results. Specifically, when the stochastic noise term in the stochastic MASs (1) vanishes, the SARE (9) degenerates into a standard algebraic Riccati equation $A^{T}P+PA-PBB^{T}P+\mathcal{Q}=0$ ($\mathcal{Q}=I_{n}$) applicable to deterministic MASs. This special case aligns perfectly with the results in [35], demonstrating the theoretical consistency between our approach and previous works.
Remark IV.4.
Theorems IV.1 and IV.2 established a unified and flexible framework for the design of fully distributed protocol in stochastic environments under directed topology. The core of the design lies in the auxiliary time-varying gain function $\Sigma_{i}(t)$, whose structure (parameterized by $k_{1}$, $k_{2}$, $\mu$, and the time-varying term $c_{i}(t)$, $\sigma_{i}(t)$) introduces significant generality in protocol design.
This generality enables our adaptive protocol to encompass several notable consensus designs for deterministic MASs under directed topologies as special cases, demonstrating broader applicability. Specifically:
  • The feedback structure proposed in [35] emerges as a special case of Remark IV.1 with $\mu=3$ and $k_{1}=k_{2}=1$.
  • The adaptive gain design presented in [14] can be obtained from Theorem IV.2 by setting $k_{1}=k_{2}=1$.
Therefore, the protocols developed in [35] and [14] can be viewed as particular parameterizations within our proposed framework (2). This generalization not only unifies existing approaches but also provides enhanced flexibility for performance tuning and robustness improvement in the presence of complex stochastic dynamics.

### 31 Mar 2026 v2

@@ IV-A Fully distributed consensus control for stochastic MASs with the directed topology
$\begin{split}\int_{0}^{\infty}\mathrm{E}\eta_{1}\varepsilon\xi^{T}(s)[\tilde{S}(\tilde{C}(s)\tilde{\Sigma}(s))\otimes P]\xi(s)ds<\infty.\end{split}$
Combining the above inequality with Barbalat’s lemma [36], we have $\lim_{t\to\infty}\mathrm{E}\xi^{T}(t)[\eta_{1}\varepsilon\tilde{S}(\tilde{C}(t)\tilde{\Sigma}(t))\otimes P]\xi(t)=0$. Note that $\tilde{C}(t)\tilde{\Sigma}(t)\geq\tilde{C}(0)>I_{N}$ and $\eta_{1},\varepsilon>0$. Therefore, we have $\lim_{t\to\infty}\mathrm{E}\|\xi(t)\|^{2}=0$. Note that $\xi(t)=(\mathcal{L}\otimes I_{n})\vartheta(t)$ and $\lim_{|\xi|\to\infty}\inf_{t\geq 0}V(\xi,t)=\infty$. Using the stochastic LaSalle lemma [37], we can get $\lim_{t\to\infty}\xi(t)=0$ a.s. According to the definition of $\xi(t)$, the m.s. and a.s. consensus can be solved. ∎
@@
In fact, the function $\Sigma_{i}$ admits other design forms. Especially, we can choose $\Sigma_{i}(t)=k_{1}(k_{2}+\sigma_{i}(t))^{\mu}$ with $\mu>1,k_{1},k_{2}\geq 1$, $c_{i}(t)\geq 1$, $\Sigma_{i}(t)\geq 1$. By selecting the Lyapunov function $\tilde{V}(\xi(t),t)$ $V(\xi(t),t)$, it can be deduced that $\mathfrak{L}\tilde{V}(\xi(t),t)=\sum_{i=1}^{M}r_{i}[\eta_{1}^{l}\Sigma_{i}(t)\sigma_{i}(t)+2\eta_{2}^{l}c_{i}(t)-2\eta_{2}^{l}\psi_{i}]\dot{c}_{i}(t)+\eta_{1}^{l}\sum_{i=1}^{M}r_{i}c_{i}(t)\Sigma_{i}(t)\mathfrak{L}\sigma_{i}(t)+\sum_{i=M+1}^{N}s_{i-M}[\eta_{1}^{f}\Sigma_{i}(t)\sigma_{i}(t)+2\eta_{2}^{f}c_{i}(t)-2\eta_{2}^{f}\psi_{i}]\dot{c}_{i}(t)+\eta_{1}^{f}\sum_{i=M+1}^{N}s_{i-M}c_{i}(t)\Sigma_{i}(t)\mathfrak{L}\sigma_{i}(t)$. By the similar steps of the above proof of Theorem IV.1, the m.s. and a.s. consensus of (1) can be solved by the protocol (2) with $\Sigma_{i}(t)=k_{1}(k_{2}+\sigma_{i}(t))^{\mu}$, for $\mu>1,k_{1},k_{2}\geq 1$ and $c_{i}(0)\geq 1$, $\Sigma_{i}(0)\geq 1$. In particular, the feedback control structure proposed in [35] is recovered as a special case of this protocol by setting $\mu=3$ and $k_{1}=k_{2}=1$.
@@
Note that the Lyapunov functional $V(\xi(t),t)$ in Theorem IV.1 satisfies $V(\xi(t),t)\leq\sum_{i=1}^{M}r_{i}[\eta_{1}^{l}k_{1}(2k_{2}c_{i}(t)+\sigma_{i}(t))\sigma_{i}(t)+\eta_{2}^{l}(c_{i}(t)-\psi_{i})^{2}]+\sum_{i=M+1}^{N}s_{i-M}[\eta_{1}^{f}k_{1}(2k_{2}c_{i}(t)+\sigma_{i}(t))\sigma_{i}(t)+\eta_{2}^{f}(c_{i}(t)-\psi_{i})^{2}]$ with $\mu=1$. Then, we choose Lyapunov functional
$\begin{split}V_{1}&(\xi(t),t)\\ =&\sum_{i=1}^{M}\!r_{i}\!\Big[\eta_{1}^{l}k_{1}(2k_{2}c_{i}(t)\!+\!\sigma_{i}(t))\sigma_{i}(t)\!+\!\eta_{2}^{l}(c_{i}(t)\!-\!\psi_{i})^{2}\Big]\\ &\!\!+\!\!\!\!\sum_{i=M+1}^{N}\!\!\!\!\!s_{i-M}\!\Big[\eta_{1}^{f}k_{1}(2k_{2}c_{i}(t)\!+\!\sigma_{i}(t))\sigma_{i}(t)\!+\!\eta_{2}^{f}(c_{i}(t)\!-\!\psi_{i})^{2}\Big],\end{split}$
where $\eta_{1}^{l},\eta_{2}^{l},\eta_{1}^{f},\eta_{2}^{f}>0$, $\psi_{i}$ are positive constants to be determined. According to the functional $\mathrm{It\hat{o}}$ formula, it can be deduced that $dV_{1}(\xi(t),t)$ has the following form
$\begin{split}dV_{\!1}&(\xi(t),t)\!=\!\mathfrak{L}V_{\!1}(\xi(t),t)dt\!+\!\frac{\partial V_{1}(\xi(t),t)}{\partial\xi}\!(I_{N}\!\otimes\!C)\xi(t)dw(t),\end{split}$
with
$\begin{split}\mathfrak{L}&V_{1}(\xi(t),t)\\ =&\sum_{i=1}^{M}\!r_{i}\big[(\eta_{k_{2}}^{l}\sigma_{i}(t)\!+\!2\eta_{2}^{l}c_{i}(t)\!-\!2\eta_{2}^{l}\psi_{i})\dot{c}_{i}(t)\big]\\ &+\sum_{i=1}^{M}\!r_{i}\big[\big(\eta_{k_{2}}^{l}c_{i}(t)\!+\!\eta_{k_{1}}^{l}\sigma_{i}(t)\big)\mathfrak{L}\sigma_{i}(t)\big]\\ &+\sum_{i=M+1}^{N}\!s_{i-M}\big[(\eta_{k_{2}}^{f}\sigma_{i}(t)\!+\!2\eta_{2}^{f}c_{i}(t)\!-\!2\eta_{2}^{f}\psi_{i})\dot{c}_{i}(t)\big]\\ &+\sum_{i=M+1}^{N}\!s_{i-M}\big[\big(\eta_{k_{2}}^{f}c_{i}(t)\!+\!\eta_{k_{1}}^{f}\sigma_{i}(t)\big)\mathfrak{L}\sigma_{i}(t)\big],\end{split}$
with $\eta_{k_{1}}^{l}=2\eta_{1}^{l}k_{1}$, $\eta_{k_{2}}^{l}=2\eta_{1}^{l}k_{1}k_{2}$, $\eta_{k_{1}}^{f}=2\eta_{1}^{f}k_{1}$, and $\eta_{k_{2}}^{f}=2\eta_{1}^{f}k_{1}k_{2}$. Denote $\tilde{\psi}=\mathrm{diag}\{\psi_{1},...,\psi_{N}\}$. According to Lemma IV.1, $\tilde{\psi}$ can be partitioned into $\tilde{\psi}_{l}=\mathrm{diag}\{\psi_{1},...,\psi_{M}\}$ and $\tilde{\psi}_{f}=\mathrm{diag}\{\psi_{M+1},...,\psi_{N}\}$. Let $\mathcal{K}=-B^{T}P$ and $\Gamma=PBB^{T}P$. Note that $\sum_{i=1}^{M}\sigma_{i}(t)=\xi_{l}^{T}(t)(I_{M}\otimes P)\xi_{l}(t)$ and $\xi_{l}(t)=(\mathcal{L}_{11}\otimes I_{n})\vartheta_{l}(t)$. By the definition of $\sigma_{i}(t)$, (24), (25) and using Lemma IV.1 and IV.2, we have
@@
Denote $\Theta^{l}_{t}=\eta_{k_{2}}^{l}\tilde{C}_{l}(t)+\eta_{k_{1}}^{l}\tilde{\sigma}_{l}(t)$ and $\Theta^{f}_{t}=\eta_{k_{2}}^{f}\tilde{C}_{f}(t)+\eta_{k_{1}}^{f}\!\tilde{\sigma}_{f}(t)$. Substituting the above inequalities into $\mathfrak{L}V_{1}(\xi(t),t)$ yields
$\begin{split}\mathfrak{L}&V_{1}(\xi(t),t)\\ \leq&\xi_{l}^{T}\!(t)[R\Theta^{l}_{t}\!\otimes\!(A^{T}\!P\!+\!PA\!+\!C^{T}\!PC)]\xi_{l}(t)\\ &+\xi_{f}^{T}\!(t)[S\Theta^{f}_{t}\!\otimes\!(A^{T}\!P\!+\!PA\!+\!C^{T}\!PC)]\xi_{f}(t)\\ &-2\tilde{\eta}_{l}\xi_{l}^{T}\!(t)\big[(k_{1}k_{2}\tilde{C}_{l}(t)\!+\!k_{1}\tilde{\sigma}_{l}(t))^{2}\!\otimes\!\Gamma\big]\xi_{l}(t)\\ &-2\tilde{\eta}_{f}\xi_{f}^{T}(t)\big[(k_{1}k_{2}\tilde{C}_{f}(t)+k_{1}\tilde{\sigma}_{f}(t))^{2}\otimes\Gamma\big]\xi_{f}(t)\\ &+\xi_{l}^{T}(t)[R(\eta_{k_{2}}^{l}\tilde{\sigma}_{l}(t)\!+\!2\eta_{2}^{l}\tilde{C}_{l}(t)\!-\!2\eta_{2}^{l}\tilde{\psi}_{l})\!\otimes\!\Gamma]\xi_{l}(t)\\ &+\xi_{f}^{T}(t)[S(\eta_{k_{2}}^{f}\tilde{\sigma}_{f}(t)\!+\!2\eta_{2}^{f}\tilde{C}_{f}(t)\!-\!2\eta_{2}^{f}\tilde{\psi}_{f})\otimes\Gamma]\xi_{l}(t),\end{split}$
@@
$\begin{split}\mathfrak{L}&V_{1}(\xi(t),t)\\ \leq&\xi_{l}^{T}\!(t)[R\Theta^{l}_{t}\!\otimes\!(A^{T}\!P\!+\!PA\!+\!C^{T}\!PC\!+\!k_{2}\Gamma)]\xi_{l}(t)\\ &\!\!\!\!+\!\xi_{f}^{T}\!(t)[S\Theta^{f}_{t}\!\otimes\!(A^{T}\!P\!+\!PA\!+\!C^{T}\!PC\!+\!k_{2}\Gamma)]\xi_{f}(t)\\ &\!\!\!\!-\!\xi_{l}^{T}\!(t)\big[(2\tilde{\eta}_{l}(k_{1}k_{2}\tilde{C}_{l}(t)\!+\!k_{1}\tilde{\sigma}_{l}(t))^{2}\!+\!2\eta_{2}^{l}\tilde{\psi}_{l}R)\!\otimes\!\Gamma\big]\xi_{l}(t)\\ &\!\!\!\!-\!\xi_{f}^{T}(t)\big[(2\tilde{\eta}_{f}(k_{1}k_{2}\tilde{C}_{f}(t)\!+\!k_{1}\tilde{\sigma}_{f}(t))^{2}\!+\!2\eta_{2}^{f}\tilde{\psi}_{f}S)\otimes\Gamma\big]\xi_{f}(t)\\ \leq&\xi_{l}^{T}\!(t)[R\Theta^{l}_{t}\!\otimes\!(A^{T}\!P\!+\!PA\!+\!C^{T}\!PC\!+\!k_{2}\Gamma)]\xi_{l}(t)\\ &\!\!\!\!+\!\xi_{f}^{T}\!(t)[S\Theta^{f}_{t}\!\otimes\!(A^{T}\!P\!+\!PA\!+\!C^{T}\!PC\!+\!k_{2}\Gamma)]\xi_{f}(t)\\ &\!\!\!\!-\!2\xi_{l}^{T}\!(t)\Big[\sqrt{\tilde{\eta}_{l}\eta_{2}^{l}\tilde{\psi}_{l}R}\Big(k_{1}k_{2}\tilde{C}_{l}(t)\!+\!k_{1}\tilde{\sigma}_{l}(t)\Big)\!\otimes\!\Gamma\big]\xi_{l}(t)\\ &\!\!\!\!-\!2\xi_{f}^{T}(t)\big[\sqrt{\tilde{\eta}_{f}\eta_{2}^{f}\tilde{\psi}_{f}S}\Big(k_{1}k_{2}\tilde{C}_{f}(t)+k_{1}\tilde{\sigma}_{f}(t)\Big)\otimes\Gamma\Big]\xi_{f}(t).\end{split}$
@@
$\begin{split}\mathfrak{L}V_{1}(\xi(t),t)\leq&-\!\xi_{l}^{T}\!(t)[R\Theta^{l}_{t}\otimes I_{n}]\xi_{l}(t)\!-\!\xi_{f}^{T}\!(t)[S\Theta^{f}_{t}\!\otimes\!I_{n}]\xi_{f}(t)\\ \leq&-\!\varepsilon\xi_{l}^{T}\!(t)\Big[R\Big(\eta_{k_{2}}^{l}\tilde{C}_{l}(t)\!+\!\frac{\eta_{k_{1}}^{l}}{2}\tilde{\sigma}_{l}(t)\Big)\otimes P\Big]\xi_{l}(t)\\ &-\!\varepsilon\xi_{f}^{T}\!(t)\Big[S\Big(\eta_{k_{2}}^{f}\tilde{C}_{f}(t)\!+\!\frac{\eta_{k_{1}}^{f}}{2}\tilde{\sigma}_{f}(t)\Big)\otimes P\Big]\xi_{f}(t)\\ \leq&-\varepsilon V_{1}(\xi(t),t)+\varphi(t),\end{split}$
@@
Theorems IV.1 and IV.2 established a unified and flexible framework for the design of fully distributed protocol in stochastic environments under directed topology. In contrast to the existing adaptive protocols [24, 26], which are designed for stochastic MASs under undirected graphs, the proposed protocols address the case of directed graphs and provide a fully distributed control framework that can be applied to solve both the m.s. and a.s. consensus under directed graphs. To this end, the adaptive strategy (2) incorporates two time-varying gains, $c_{i}(t)$ and $\Sigma_{i}(t)$, each serving a distinct role in facilitating stochastic consensus: $c_{i}(t)$ is employed to estimate global information, while $\Sigma_{i}(t)$ accounts for the asymmetrical interactions inherent in directed graphs.
@@
In fact, the structure of the auxiliary time-varying gain function $\Sigma_{i}(t)$, parameterized by $k_{1}$, $k_{2}$, $\mu$, and the time-varying terms $c_{i}(t)$, $\sigma_{i}(t)$, introduces generality into the protocol design. This generality allows our adaptive protocol to encompass several representative consensus designs for deterministic multi-agent systems under directed topologies as special cases, thereby demonstrating broader applicability. For instance, the adaptive gain design in [14] can be obtained from Theorem IV.2 by setting $k_{1}=k_{2}=1$, which can be regarded as specific parameterizations within our proposed framework (2).